# Self-Supervised Skeleton Action Representation Learning: A Benchmark and Beyond

### v1

@@ 4.6 Benchmark on Different Downstream Tasks
$\bullet$ Unsupervised Learning Setting follows the linear evaluation protocol, adding a trainable linear classifier after the fixed pre-trained encoder. As shown in Table LABEL:tab:main_comp, PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols. Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
$\bullet$ Semi-supervised Learning Setting fine-tunes the whole model with only a portion of labeled data. This reflects the effectiveness in terms of avoiding over-fitting problem of the pretrained model. In addition to fine-tuning the whole model, we find the linear evaluation protocol tends to yield a better performance, especially for the small ratio of labeled data. As shown in Table II, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data. In contrast, previous works based on reconstruction [7] or contrastive learning [62, 10] solely can not achieve satisfactory results.
TABLE III: The results of action retrieval with joint stream.
@@
TABLE IV: The results of occluded action recognition. $\Delta_{\downarrow}$ represents the average performance degradation compared to that without occlusion.
@@
Spatial Occ. (%) Temporal Occ. (%)
xsub xview $\Delta_{\downarrow}$ xsub xview $\Delta_{\downarrow}$
MoCo-GRU [49] 64.8 72.6 12.4 68.8 74.8 9.3
ISC [62] 62.8 70.6 14.1 68.9 76.8 7.9
CRRL [15] 56.8 61.4 11.6 61.0 66.2 7.1
CMD [10] 67.1 72.7 13.3 72.7 79.5 7.1
HiCo [11] 66.4 72.1 15.4 71.2 76.7 10.6
PCM3 [71] 80.8 87.0 3.3 77.6 86.1 5.4
PCM3++ 81.8 88.0 3.0 79.9 85.8 5.0
2) Skeleton-based Action Retrieval. Following previous work [14], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space. The precision is reported as accuracy in Table III. Our method achieves a competitive performance compared with other latest methods. This indicates the highly distinguishable representations learned by our method. Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [7] and MAMP [48].
[…]
@@
TABLE VI: Comparison of few shot learning for skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot task.
@@
As presented in Table IV, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion. Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity. However, due to the lack of modeling learning of occluded data, other methods are with poor robustness for occlusion.
4) Skeleton-based Action Detection. We follow the previous works [96, 28], and evaluate the short-term motion modeling capacity by action detection task. We only train the attached linear classifier (fully-connected layer) after the pre-trained encoder to predict frame-level categories and formulate the final proposal. The encoder is pre-trained on NTU 60 dataset and then we transfer the learned representations to the untrimmed PKUMMD Part I dataset. The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table V. Note that we provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline. First, it can be found that the SSL pre-training can significantly improve the performance compared with the baseline model. Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [10], HiCo [11], and LAC [99], which ignore the clip-level representation modeling.
TABLE VII: Ablation study on the clip-level contrastive learning, the prompted training, and the post-distillation.
Method Recognition (%) Detection (%)
@@
TABLE VIII: Ablation study on the representation learning at different levels. Multi-Task denotes their simple combination of contrastive learning and masked prediction. $J,C,S$ represents the features of joint, clip, and sequence levels.
Method Representation Recognition Retrieval Recognition w Occlusion Detection Few Shot Learning
Acc. (%) Acc. (%) Acc. (%) mAP@0.1 (%) Acc. (%)
Masked Prediction J 14.7 64.4 11.7 36.1 43.5
Contrastive Learning S 87.3 84.5 76.6 72.5 61.4
Multi-Task J + S 87.5 85.1 77.4 71.9 60.2
PCM3 J + S 90.4 88.8 87.0 73.3 65.2
PCM3++ J + C + S 91.0 89.4 88.0 75.5 66.7
5) Unsupervised Few-shot Learning. In this task, we evaluate the performance of SSL as few-shot learners follow the previous work [101]. Specifically, the model is first pre-trained on NTU 60 dataset. Then, a simple classifier, e.g., Support Vector Machine (SVM) in our implementation, is fitted on the output features by the pre-trained encoder of the support data set. Finally, the adapted classifier along with the encoder is utilized to infer the query samples. We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected. which are not seen in pre-training, as the support and query sets following [102]. In Table VI, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training specifically for unseen categories.
6) Compared with the Baseline Method. We evaluate and present the comparison results of our baseline model, PCM3 [71], on various downstream tasks as shown in Table LABEL:tab:main_comp-VI. As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy. These methodological advancements consistently improve the model’s representation quality and robustness on various downstream tasks, especially the semi-supervised action recognition and action detection performance.
To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding. Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table III and Table VI, but not equal to greater robustness to occlusion. Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling. We design a novel asymmetric clip-level contrastive learning pretext task and effectively boost the detection performance. Remarkably, our proposed method achieves promising results on these tasks compared with other methods that ignore the versatile representation learning of different granularity.
@@ 4.7 Analysis of PCM3++
4.7.1 Ablation Study of Relevant Components
1) Analysis of clip contrastive learning. We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table VII. Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity.
2) Analysis of the versatile representation learning. As shown in Table VIII, we analyze the performance under different downstream tasks when adopting single paradigm or combining them. First, as discussed before, only employing the masked skeleton modeling cannot obtain a distinguishable representation space for GRU, leading to poor results in linear discrimination task. On the other hand, contrastive learning can achieve decent results on different downstream tasks by model the high-level semantics. However, it only learns the sequence-level representations, leaving the generalization capacity under-explored. Therefore, to achieve more representative features, we can combine different paradigms. However, simply combining them only shows mediocre improvement because it neglects the connection between the two tasks. In contrast, PCM3 [71] utilizes the synergy between the two tasks, and significantly improve the performance and generalization capacity. In addition, we integrate the clip-level representations and achieve further improvement.
3) Effect of the post-distillation design. As shown in Table VII, the post-distillation training strategy can bring further improvement slightly for the recognition task. It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
TABLE IX: Action recognition results using different backbones with joint stream.
Method Backbone xsub (%) xview (%)
@@ 4.7.1 Ablation Study of Relevant Components
4.7.2 More Analysis Results
1) Results of different backbones. We give more results with different backbones in Table IX. Specifically, for GCN and Transformer, we adopt ST-GCN [4] and DSTA-Net [98] following previous works [67, 12]. As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods. Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is ,however, not the focus of this paper, and we leave it for future work. We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
2) Complexity Analysis. We give an analysis of space and computational complexities of our method for pre-training in Table X. As we can see, compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity. For the space complexity, the main additional cost is the reconstruction decoder $dec(\cdot)$, which takes up about 4M of space. For the computational complexity, the encoding process of different positive samples contributes most of the computational overhead.
TABLE X: FLOPs and Params results of different models.
Models Params $\downarrow$ FLOPs $\downarrow$ Accuracy
@@ 5 Conclusions and Future Directions
This paper presents a comprehensive survey on self-supervised skeleton-based action representation learning, where different literature is organized following the taxonomy of context-based, generative learning, and contrastive learning approaches. Then, a detailed benchmark and insightful discussions are provided, including the datasets, model backbones and pretext tasks. After reviewing existing works, we technically propose a novel and effective framework, for versatile skeleton-based action representation learning which is less explored before as a challenging topic. Revisiting the combination of contrastive learning and masked skeleton modeling paradigms, our method can achieve representation learning of different granularity by fully utilizing the novel spatial-temporal motion patterns. Extensive experiments on five downstream tasks demonstrate our superior performance and generalization capacity.
@@
  • Long-Term Motion Understanding. Existing methods utilize the trimmed video clip as training data, i.e., a video sample only contains one motion. This limits the long-term temporal reasoning capacity, especially for long videos containing multiple actions, e.g., Long-Term Action Anticipation, which should be paid more efforts.
  • Multi-Modal Learning. The exploration of multi-modal pre-training including human skeleton data is still insufficient. For example, as a useful complement to the skeleton data, RGB data can provide additional background knowledge that can boost action representation learning. Meanwhile, although valuable efforts have been made on the skeleton-text alignment [103, 104], stronger, larger-scale pre-training models are still urgently needed.
  • Versatile Representation Learning. As discussed in previous sections, more downstream tasks should be involved to fully explore and exploit the generalization capacity of SSL models. In addition to combining different pre-training paradigms analogous to this work, new frameworks can also be explored, e.g., Diffusion model [105], which is promising to handle both the skeleton-based generative and discriminative downstream tasks.
  • Towards Skeleton in the Wild. Existing methods are mostly evaluated in simplified and controlled environments and can suffer from serious noise caused by occlusion and view variation when deployed on a more diverse outdoor scenario. As an effective technique to improve model generalization, SSL is promising to boost skeleton representation robustness in the wild.
In summary, many abundant practices have emerged in skeleton-based SSL literature, and more valuable endeavours are expected to improve current SSL works and explore new directions.

### v2

@@ 4.6 Benchmark on Different Downstream Tasks
Firstly, the unsupervised learning results using linear evaluation protocol is shown in Table LABEL:tab:main_comp, and PCM3++ achieves better or competitive results compared with other state-of-the-art methods on different evaluation protocols. Meanwhile, the results on PKUMMD dataset can be found in supplementary material, which also indicates the strong generalization capacity on the noisy data.
In addition, we also report the semi-supervised setting results, which fine-tunes the whole model/classifier with only a portion of labeled data. As shown in Table 2, PCM3++ obtains better performance than both the semi- and self-supervised methods with different proportions of training data, and only fine-tuning the classifier tends to be better with extremely little data. In contrast, previous works based on reconstruction [100] or contrastive learning [72, 52] solely can not achieve satisfactory results.
Table 3: The results of action retrieval with joint stream.
@@
Table 4: The results of occluded action recognition.
@@
Spatial Occ. Temporal Occ.
xsub xview xsub xview
MoCo-GRU [26] 64.8 72.6 68.8 74.8
ISC [72] 62.8 70.6 68.9 76.8
CRRL [78] 56.8 61.4 61.0 66.2
CMD [52] 67.1 72.7 72.7 79.5
HiCo [16] 66.4 72.1 71.2 76.7
PCM3 [98] 80.8 87.0 77.6 86.1
PCM3++ 81.8 88.0 79.9 85.8
2) Skeleton-based Action Retrieval. Following previous work [68], a K-nearest neighbors (KNN) classifier ($k$=1) is adopted to retrieve the nearest training sample for each testing data in the representation space. The precision is reported as accuracy in Table 3. Our method achieves a competitive performance compared with other latest methods. This indicates the highly distinguishable representations learned by our method. Meanwhile, it is found that contrastive learning based methods usually perform better than reconstruction-based methods, e.g., LongT GAN [100] and MAMP [51].
Table 5: Action detection results on PKUMMD Part I benchmark under linear evaluation protocol.
@@
Table 6: Comparison of few shot skeleton-based action recognition. ($k$, $n$) represents $k$-way $n$-shot.
@@
As presented in Table 4, our method can capture underlying structures in the distorted data owing to the masked contrastive learning, and is good at dealing with the spatial occlusion. Meanwhile, this capacity extends well for temporal occlusion, demonstrating a desirable generalization capacity. However, due to the lack of modeling learning of occluded data, other methods are with poor robustness for occlusion.
4) Skeleton-based Action Detection. We follow the previous works [49, 43], and evaluate the short-term motion modeling capacity by action detection task. We train the attached linear classifier to predict frame-level categories to produce the final proposal. The encoder is pre-trained on NTU 60 dataset and then we fix it, transferring the learned representations to untrimmed PKUMMD Part I dataset. The mean average precision (mAP) of different actions is adopted as metric with different temporal overlapping ratios (tIoU) in Table 5. We provide a fully fine-tuned model (train both the encoder and classifier) without pre-training as the baseline. First, it can be found the SSL pre-training can significantly improve the performance compared with the baseline model. Besides, our method integrates the clip-level representation contrastive learning and achieves the best scores compared with latest CMD [52], HiCo [16], and LAC [88], which ignore the clip-level representation modeling.
Table 7: Ablation study on the clip-level contrastive learning and the post-distillation.
Method Recognition Detection
@@
5) Unsupervised Few-shot Learning. In this task, we evaluate the performance of SSL as few-shot learners follow the previous work [48]. Specifically, the model is first pre-trained on NTU 60 dataset. Then, a simple classifier, e.g., Support Vector Machine (SVM) in our implementation, is fitted on the output features by the pre-trained encoder of the support data set. Finally, the adapted classifier along with the encoder is utilized to infer the query samples. We select 20 new categories in NTU 120111Specifically, the classes (index from 0) 60, 61, 66, 69, 72, 78, 79, 80, 84, 90, 91, 95, 96, 98, 99, 100, 102, 106, 108, 111, 113 114, 115 are selected. which are not seen in pre-training, as the support and query sets following [46]. In Table 6, our method generalizes well on the unseen categories and surpasses the supervised baseline and other methods notably, although there is no training for unseen categories.
6) Compared with the Baseline Method. We evaluate and present the comparison results of our baseline model, PCM3 [98], on various downstream tasks as shown in Table LABEL:tab:main_comp-6. As our new improvement, we integrate an effective clip-level contrastive learning scheme and present a novel post-distillation training strategy. These methodological advancements consistently improve the model’s representation quality and robustness on various downstream tasks, especially the semi-supervised action recognition and action detection performance.
To sum up, self-supervised learning provides a strong solution to boost various downstream tasks for versatile skeleton-based action understanding. Generally, a higher recognition performance roughly indicates a better representation space for most downstream tasks, as found in Table 3 and Table 6, but not equal to greater robustness to occlusion. Meanwhile, action detection is another crucial task for action analysis, which mainly focus on the short-term motion modeling. We design a novel asymmetric clip-level contrastive learning pretext task and effectively boost the detection performance. Remarkably, our proposed method achieves promising results on these tasks compared with other methods that ignore the versatile representation learning of different granularity.
@@ 4.7 Analysis of PCM3++
1) Analysis of clip contrastive learning. We add the temporal clips as the asymmetric positive sample of the anchor sequence, which further improves the representation quality as shown in Table 7. Remarkably, the clip-level contrastive learning significantly boost the action detection performance by promoting short-term modeling capacity. Our designs boost versatile representation learning of different granularity and the effect on more downstream tasks can be found in Supp.
2) Effect of the post-distillation design. As shown in Table 7, the post-distillation training strategy can bring further improvement slightly for the recognition task. It alleviates the false negative problem by removing the one-hot pseudo label in the contrastive learning, achieving a more compact representation space.
Table 8: Results of different backbones.
Method Backbone xsub xview
@@
Table 9: FLOPs and Params results.
[…]
@@
3) Results of different backbones. We give more results with different backbones in Table 8. Specifically, for GCN and Transformer, we adopt ST-GCN [86] and DSTA-Net [64] following previous works [41, 97]. As we can see, our method shows good generalization capacity across different model backbones, and performs better or on-par compared with the latest SSL methods. Additional improvement can be potentially achieved by further searching for hyper-parameters on different backbones, which is, however, not the focus of this paper, and we leave it for future work. We finally adopt GRU model in implementation for the more friendly memory usage of GPU and the higher performance.
4) Complexity Analysis. We give an analysis of space and computational complexities of our method for pre-training in Table 9. Compared with other GRU-based and Transformer-based methods, our method achieves a significant performance improvement with an acceptable cost of the complexity. For the space complexity, the main extra cost is the reconstruction decoder (4M of space). For the computational complexity, the encoding process of different positive samples contributes most of the overhead.
@@ 5 Conclusions and Future Directions
This paper presents a comprehensive survey on skeleton-based action representation SSL, where different literature is organized following the taxonomy of context-based, generative learning, and contrastive learning approaches. After reviewing existing works, we technically propose a novel and effective framework, for versatile skeleton-based action representation learning which is less explored before as a challenging topic. Finally, a detailed benchmark and insightful discussions are provided, including datasets, model backbones and pretext tasks. Meanwhile, we demonstrate the superior generalization performance of our method under different downstream tasks.
@@
  • Long-Term Motion Understanding. Existing methods utilize the trimmed video clip, only containing one motion, as training data. This limits the long-term temporal reasoning capacity containing multiple actions, e.g., Long-Term Action Anticipation Task.
  • Multi-Modal Learning. The exploration of multi-modal pre-training including human skeleton data is still insufficient, e.g., text, infrared data, and RGB images. For example, as a useful complement to skeleton, RGB data can provide additional background knowledge that can boost action representation learning.
  • Towards Skeleton in the Wild. Existing methods are evaluated in simplified and controlled environments and can suffer from serious noise, e.g., occlusion and view variation when deployed on the outdoor scenario. On this point, SSL is promising to boost skeleton representation robustness in the wild.
  • Versatile Representation Learning. As discussed in previous sections, most works only focus on the recognition task, leaving the generation capacity across different tasks of SSL models under-explored. Meanwhile, new frameworks can also be explored, e.g., Diffusion model [27], which is promising to handle both generative and discriminative downstream tasks.
Declarations
  • Competing interests. This work was supported in part by the National Natural Science Foundation of China under Grant No.62172020, and in part by the Key Laboratory of Science, Technology and Standard in Press Industry (Key Laboratory of Intelligent Press Media Technology).
  • Data availability. The adopted NTU-RGB+D [62, 44] and PKU-MMD [43] dataset are popular public benchmarks in skeleton-based action understanding. The code for data processing can be found in https://github.com/JHang2020/PCM3.